OpenStudy (anonymous):

What are the zeroes of f(x) = x2 - 6x + 8?

OpenStudy (anonymous):

x = -4, 2 x = -4, -2 x = 4, 2 x = 4, -2

OpenStudy (solomonzelman):

Set the function equal zero? (I mean, x^2 - 6x + 8 = 0 ) Factor . Solve for x. let me know if you still need help.

OpenStudy (anonymous):

A

OpenStudy (solomonzelman):

x^2 - 6x + 8 = 0 (x - 4) (x - 2) = 0 know which option do you choose?

OpenStudy (solomonzelman):

meant to say "now" , not "know"

OpenStudy (anonymous):

B

OpenStudy (solomonzelman):

Are you guessing ?

OpenStudy (anonymous):

yes and no

OpenStudy (anonymous):

i think it b